OpenStudy (anonymous):

Write a scenario to reflect the following equation. 2x+4=12

OpenStudy (lin.ivory):

2x=12-4 2x=8 2x/2=8/2 x=4

OpenStudy (anonymous):

i have 2 glasses of beer, and another 4 ounces of beer the total number of ounces of beer i have is 12 how much beer is in each glass?